# Jingjing Sun, Pengyi Shi, Jim Dai

### v3

@@ 4.1.1 Randomized Atomic Action
In this paper, we use randomized policies to prescribe the atomic actions. We will first outline the specifics of this randomization and then discuss its advantages over conventional deterministic policies in Section 4.1.2.
For a given atomic state $s^{n}$, we define a probability distribution over all possible atomic actions for each customer class. We denote this distribution for a class $i$ customer, given $s^{n}$, as $\kappa(\cdot|s^{n},i)$, and the set of distributions as $\kappa=\{\kappa(\cdot|s^{n},i),i\in\mathcal{J}\}$. Consequently, starting from the initial state $s^{1}=s$ (the system-level pre-action state) and following a given order $\boldsymbol{\sigma}(s)$, the probability of selecting a sequence of atomic actions $\textbf{a}=\{a^{1},\dots,a^{q}\}$ is given by $\prod_{n=1}^{q}\kappa(a^{n}|s^{n},\sigma^{n}(s))$, where $(s^{1},\dots,s^{q})$ denotes the corresponding atomic state sequence generated within the atomic decision process, as specified in Equation (4).
@@
$\pi(f|s)=$ $\sum_{\textbf{a}}\prod_{n=1}^{q}\kappa(a^{n}|s^{n},\sigma^{n}(s))\cdot\mathbbm{1}\Big{(}\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i)=f_{i,j},\forall i,j\Big{)}.$ (5)
To simplify notation and align with the broader PPO algorithm literature, we use the system-level policy $\pi=\{\pi(f|s)\}$ in the rest of the paper when we refer to a policy. However, it is important to clarify that this system-level policy is driven by the underlying atomic-level probability distributions. In other words, policy $\pi$ is the outcome of the sequence of atomic actions governed by the distributions $\kappa$. Our goal is to optimize $\kappa$ in such a way that the induced system-level policy $\pi$ achieves the performance objective in (3). Moreover, in Section 5.3, we will introduce a “batching” setup that simplifies the calculation of (5) and eliminates the dependence on the specific processing order $\boldsymbol{\sigma}$. This will establish a more direct connection between the system-level policy $\pi$ and the atomic-level distributions $\kappa$. We discuss how the randomized policy can be practically applied in decision-making in Section 6.5.
@@ 4.1.2 Policy Parameterization
where $g_{i,j}(s|\theta)$ is one of the $J\times J$ output neurons of a dense NN with input $s$ and parameters $\theta$. In the simplest setting, we can use a NN with no hidden layers, i.e.,
$g_{i,j}\big{(}s|\theta=(w_{i,j},b_{i,j},i,j\in\mathcal{J})\big{)}=w_{i,j}\cdot s+b_{i,j},$
which is equivalent to a multi-class logistic regression model. For further details on NN parameterization, see Section 5.1.
Note that Equation (6) could assign a positive probability to non-feasible atomic actions, such as when the selected server pool is full or cannot serve class $j$. To ensure the feasibility of the prescribed atomic action, we set $\kappa_{\theta}(a^{n}|s^{n},\sigma^{n}(s))=0$ for $a^{n}\not\in\mathcal{A}(s^{n},\sigma^{n}(s))$ (i.e., force the probability of non-feasible actions to 0), and then rescale the remaining probabilities to sum to 1. This adjustment is equivalent to resampling: if a sampled atomic action based on $\kappa_{\theta}$ is not feasible, we continue resampling until a feasible action is selected.
@@ 4.2 Finding Optimal Policy: Proximal Policy Optimization
The randomized policy framework enables the use of policy gradient methods to learn the parameter $\theta$. In this paper, we employ the state-of-the-art policy gradient method, PPO, with custom modifications tailored to our specific queueing network context. We provide an overview of the PPO framework in this section; we defer the details of our tailored modifications to Section 5.
@@
$L(\theta):=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}s\sim\mu_{\eta}\\ f\sim\pi_{\eta}(\cdot|s)\end{subarray}}\max\left\{r_{\theta,\eta}(f|s)A_{\eta}(s,f),\quad\text{clip}\left(r_{\theta,\eta}(f|s),1-\epsilon,1+\epsilon\right)A_{\eta}(s,f)\right\}.$ (9)
@@
$r_{\theta,\eta}(f|s):=\frac{\pi_{\theta}(f|s)}{\pi_{\eta}(f|s)},$ (10)
where $\pi_{\theta}(f|s)$ and $\pi_{\eta}(f|s)$ are defined as in (5), with $\kappa$ replaced by $\kappa_{\theta}$ and $\kappa_{\eta}$, respectively; and the clipped function
@@
The acceptable difference $\epsilon$ is a hyper-parameter that can be tuned for better algorithm performance.
For the theoretical guarantee, we establish Proposition 1, which extends the performance improvement guarantee of PPO from the time-stationary, long-run average cost setting in Dai and Gluzman (2022) to our more complex periodic environment. See the complete statement of Proposition 1 (alongside needed technical conditions) in Appendix 10 and its proof in Appendix 13. The key challenge in the proof lies in the periodicity of state transitions, which makes directly applying previous results inapplicable. To address this, we introduce a new framework using daily Markov chains (MCs), where one period (e.g., one day) is treated as a single transition step. This transformation allows us to redefine the objective function in a way that ensures aperiodicity and enables the application of previous theoretical framework. Proposition 1 establishes that the change in long-run average cost can be decomposed into two terms: one that dominates the improvement and another that diminishes at a faster rate. Since PPO minimizes the dominant term while keeping the policy update conservative to control the secondary term, it leads to a provable policy improvement guarantee.
@@
$\hat{L}(\theta,\mathcal{D}_{\eta}^{T},\hat{\textbf{A}}_{\eta}(\mathcal{D}_{\eta}^{T}))=\frac{1}{Tm}\sum_{t=1}^{T}\sum_{k=0}^{m-1}\max$$\left\{r_{\theta,\eta}(f(t_{k})|s(t_{k}))\hat{A}_{\eta}(s(t_{k}),f(t_{k})),\right.$ (11)
$\left.\text{clip}\left(r_{\theta,\eta}(f(t_{k})|s(t_{k})),1-\epsilon,1+\epsilon\right)\hat{A}_{\eta}(s(t_{k}),f(t_{k}))\right\},$
where the dataset $\mathcal{D}_{\eta}^{T}=\left\{\big{(}s(t_{k}),f(t_{k}),g(t_{k})\big{)}_{k=0}^{m-1}\right\}_{t=1}^{T}$ contains $T$ days of simulation data generated under policy $\pi_{\eta}$, and the vector $\hat{\textbf{A}}_{\eta}(\mathcal{D}_{\eta}^{T})=\left\{\left(\hat{A}_{\eta}(s(t_{k}),f(t_{k}))\right)_{k=0}^{m-1}\right\}_{t=1}^{T}$ consists of the estimated advantage functions for each decision epoch within the $T$ days using the collected data $\mathcal{D}_{\eta}^{T}$. Note that the constant factor $\frac{1}{Tm}$ in $\hat{L}$ can be disregarded without affecting the objective optimization. The detailed steps of the PPO algorithm are given in Algorithm 1.
@@
1:  for policy iteration $r=0,1,2,...,R-1$  do
2:     Run policy $\pi_{\eta}$ for $T$ days and collect simulation dataset $\mathcal{D}_{\eta}^{T}=\left\{\big{(}s(t_{k}),f(t_{k}),g(t_{k})\big{)}_{k=0}^{m-1}\right\}_{t=1}^{T}$.
@@
We conclude this section with two important remarks. First, our policy parameterization differs from the conventional ones used in PPO framework, which would directly parameterize a randomized policy over the entire combinatorial system-level action space. This action space has a size of $\prod_{j=1}^{J}\left(\begin{array}[]{c}q_{j}+w_{j}\\ w_{j}\end{array}\right)$, which means that a conventional policy network would require an extremely large output layer, making training highly challenging. The key element for our algorithm is the combination of atomic actions and the randomized policy. Second, performance improvement using the PPO algorithm is ensured when policy updates are conservative (see Appendix 10). The atomic action probabilities are parameterized using continuous functions that output distributions over small action spaces, allowing incremental updates that are scalable given the relatively small parameter size. This would not be feasible if we use the deterministic policy setup. We have also conducted numerical experiments that demonstrate the superior performance of our randomized approach compared to deterministic ones. The numerical results highlight two major advantages of randomized atomic actions: (i) avoiding chattering – randomized policies integrate well with PPO, which prevent policy “chattering” in discrete-action spaces; (ii) randomized actions naturally allow for exploration, which improve the speed of convergence to a “good” final policy. We discuss how to use the randomized policy in practice in Section 6.5.
5 PPO Design: Tailoring for Periodicity in MDP
@@ 5.1 Policy Network Design
Figure 2: Three policy network structures. The fully-connected network takes epoch index as part of the input and uses a single network to learn actions across all epochs. The fully-separate network trains $m$ separate networks for each of the $m$ epochs. The partially-shared network excludes the epoch index from input but uses different blocks of output neurons for actions of different epochs. The parameters of the first several layers are shared across epochs and while those of the later layers are specific to each epoch.
@@
    where $g_{i,j}$ is one of the $J\times J$ output neurons of the fully-connected network. This design is the most straightforward way to handle time-dependency – the network uses the same set of parameters to generate outputs for different epochs, which captures temporal differences by incorporating time as part of the input.
@@
Combining advantages. The fully-connected and fully-separate designs each have their pros and cons, while the partially-shared policy network combines their strengths. Specifically, the fully-connected network design uses the same set of parameters across all time epochs, naturally capturing similarities between epochs and enabling training with data from all time epochs. However, since the policy differences between epochs can be highly complex, a more sophisticated network (e.g., with additional hidden layers) may be needed to capture these nuances and hence more training data. The fully-separate network design accounts for differences between epochs with independent networks, but it would also require more time and data to train, as each network is only trained using data from its corresponding epoch. The partially-shared network design strikes a balance: shared parameters in the early layers capture commonalities between epochs for efficient training, while the separated last layer adapts to epoch-specific policy outputs. In the case study of a ten-pool system, the partially-shared design requires only a few thousand simulation data samples for training, yet outperforms the other two designs that need 100 times more data. We will provide a detailed comparison of the sample efficiency of these three designs in Section 6.4.
@@ 5.2 Policy Evaluation: Linear Value Function Approximation
In the standard PPO implementation, a separate neural network, known as the value network, is used alongside the policy network to approximate value functions. The value network is trained with targets estimated via Monte Carlo evaluation. These estimates can have high variance in long-run average settings, which require large amounts of data to achieve accurate policy evaluation. To address this, we adopt a linear value function approximation and leverage the queueing system structure to design proper basis functions. This approach mitigates variance-related inefficiencies and contrasts with the black-box nature of NN-based value estimation. Additionally, our approach extends to other operations problems with similar structural properties.
@@
Specifically, under the randomized atomic policy $\pi$, we decompose the original $J$-pool system into $J$ single-pool systems. For each decomposed pool-$j$ system, the relative value function $\hat{v}_{\pi,j}$ is solved via the corresponding post-action Poisson equation
@@
for a given post-action state $s_{j}$ (we drop the superscript $+$ here for simplicity). Here, $C_{j}q_{j}$ is the post-action holding cost with $q_{j}=\max(s_{j}-N_{j},0)$, $\hat{\gamma}_{\pi,j}$ is the average cost of the decomposed pool-$j$ system, and the cost-to-go term consists of: (a) the expected overflow cost $\mathbb{E}[\sum_{k=1}^{J}B_{j,k}\bar{f}_{j,k}]$, where $\bar{f}_{j,k}$ approximates the overflow assignments from class $j$ to other pools, and (b) the expected value function at the next post-action state $\mathbb{E}\left[\hat{v}_{\pi,j}(s’_{j})\right]$. To evaluate (a) and (b), we need to properly adjust the queueing dynamics for each decomposed pool-$j$ system by incorporating overflow assignments from other classes while excluding overflow assignments from class $j$ to other pools (treated as “diversions”). To achieve this, we approximate the probability of assigning a class-$j$ customer to pool $k$ under policy $\pi$ using $\bar{\kappa}_{j}(k|s_{j},j)$, which depends only on pool-$j$’s state information. Given the arrivals are Poisson and leveraging the Poisson thinning property, the outflows $\bar{f}_{j,k}$ and inflows $\bar{f}_{i,j}$ can be approximated as Poisson distributions with rates $\lambda_{j}\bar{\kappa}_{j}(k|s_{j},j)$ and $\lambda_{i}\bar{\kappa}_{i}(j|s_{j},i)$, respectively. This formulation allows efficient evaluation of both (a) and (b). Details on computing $\bar{\kappa}_{j}$, deriving $\hat{v}_{\pi,j}$, and formulating (13) using post-action states are provided in Appendix 11.
The feasibility of the pool-wise decomposition relies on the randomized atomic policy. The continuous nature of overflow probabilities under this policy allows us to generate good approximations of the $f_{i,j}$ values. In contrast, a deterministic policy results in discrete actions with more drastic variations across states, making it harder to approximate overflow proportions effectively. This distinction makes our new basis function, $V_{d}$, specifically suited for the PPO framework and different from the $V_{s}$ basis function proposed in Dai and Shi (2019), which was designed for deterministic, discrete-action spaces. In Section 6.4, we compare the performance of $V_{d}$ and $V_{s}$ alongside other polynomial bases. Our results demonstrate that $V_{d}$ improves sample efficiency and value function approximation accuracy, which ultimately leads to better policy performance.

### v4

@@ 4.1.1 Randomized Atomic Action
In this paper, we use randomized policies to prescribe the atomic actions. We will first outline the specifics of this randomization and then discuss its advantages over conventional deterministic policies at the end of Section 4.2.
For a given atomic state $s^{n}$, we define a probability distribution over all possible atomic actions for each customer class. We denote this distribution for a class $i$ customer, given $s^{n}$, as $\kappa(\cdot|s^{n},i)$, and the set of distributions as $\kappa=\{\kappa(\cdot|s^{n},i),i\in\mathcal{J}\}$. Consequently, starting from the initial state $s^{1}=s$ (the system-level pre-action state) and following a given order $\boldsymbol{\sigma}(s)=\{\sigma^{1}(s),\dots,\sigma^{q}(s)\}$ for $q$ waiting customers, the probability of selecting a sequence of atomic actions $\textbf{a}=\{a^{1},\dots,a^{q}\}$ is given by $\prod_{n=1}^{q}\kappa(a^{n}|s^{n},\sigma^{n}(s))$, where $(s^{1},\dots,s^{q})$ denotes the corresponding atomic state sequence generated within the atomic decision process, as specified in Equation (4).
@@
$\pi(f\mid s)=$ $\sum_{\textbf{a}}\prod_{n=1}^{q}\kappa(a^{n}|s^{n},\sigma^{n}(s))\cdot\mathbbm{1}\Big(\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i)=f_{i,j},\forall i,j\Big).$ (5)
To simplify notation and align with the broader PPO algorithm literature, we use the system-level policy $\pi=\{\pi(f\mid s)\}$ in the rest of the paper when we refer to a policy. However, it is important to clarify that this system-level policy is driven by the underlying atomic-level probability distributions. In other words, policy $\pi$ is the outcome of the sequence of atomic actions governed by the distributions $\kappa$. Our goal is to optimize $\kappa$ in such a way that the induced system-level policy $\pi$ achieves the performance objective in (3). We discuss how the randomized policy can be practically applied in decision-making in Section 6.5.
@@ 4.1.2 Policy Parameterization
where $g_{i,j}(s|\theta)$ is one of the $J\times J$ output neurons of a dense NN with input $s$ and parameters $\theta$.
Equation (6) could assign a positive probability to non-feasible atomic actions, such as when the selected server pool is full or cannot serve class $j$. To ensure the feasibility of the prescribed atomic action, we set $\kappa_{\theta}(a^{n}|s^{n},\sigma^{n}(s))=0$ for $a^{n}\not\in\mathcal{A}(s^{n},\sigma^{n}(s))$ (i.e., force the probability of non-feasible actions to 0), and then rescale the remaining probabilities to sum to 1. This adjustment is equivalent to resampling: if a sampled atomic action based on $\kappa_{\theta}$ is not feasible, we continue resampling until a feasible action is selected.
@@ 4.2 Finding Optimal Policy: Proximal Policy Optimization
We employ the state-of-the-art policy gradient method, PPO, to learn the parameter $\theta$ with custom modifications tailored to our specific queueing network context. We provide an overview of the PPO framework in this section; we defer the details of our tailored modifications to Section 5.
@@
$L(\theta):=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}s\sim\mu_{\eta}\\ f\sim\pi_{\eta}(\cdot|s)\end{subarray}}\max\left\{r_{\theta,\eta}(f\mid s)A_{\eta}(s,f),\quad\text{clip}\left(r_{\theta,\eta}(f\mid s),1-\epsilon,1+\epsilon\right)A_{\eta}(s,f)\right\}.$ (9)
@@
$r_{\theta,\eta}(f\mid s):=\frac{\pi_{\theta}(f\mid s)}{\pi_{\eta}(f\mid s)},$ (10)
where $\pi_{\theta}(f\mid s)$ and $\pi_{\eta}(f\mid s)$ are defined as in (5), with $\kappa$ replaced by $\kappa_{\theta}$ and $\kappa_{\eta}$, respectively; and the clipped function
@@
The difference $\epsilon$ is a hyper-parameter that can be tuned for better algorithm performance.
For the theoretical guarantee, we establish Proposition 1, which extends the performance improvement guarantee of PPO from the time-stationary, long-run average cost setting in Dai and Gluzman (2022) to our more complex periodic environment. See the complete statement of Proposition 1 (alongside needed technical conditions) in Appendix 10. The key challenge in the proof lies in the periodicity of state transitions, which makes directly applying previous results inapplicable. To address this, we introduce a new framework using daily Markov chains (MCs), where one period (e.g., one day) is treated as a single transition step. This transformation allows us to redefine the objective function in a way that ensures aperiodicity to use previous theoretical frameworks. Proposition 1 establishes that the change in long-run average cost can be decomposed into two terms: one that dominates the improvement and another that diminishes at a faster rate. Since PPO minimizes the dominant term while keeping the policy update conservative to control the secondary term, it leads to a provable policy improvement guarantee.
@@
$\hat{L}(\theta,\mathcal{D}_{\eta}^{T},\hat{\textbf{A}}_{\eta}(\mathcal{D}_{\eta}^{T}))=\frac{1}{Tm}\sum_{t=1}^{T}\sum_{k=0}^{m-1}\max$ $\left\{\hat{r}_{\theta,\eta}(f(t_{k})|s(t_{k}))\hat{A}_{\eta}(s(t_{k}),f(t_{k})),\right$. (11)
$\left.\text{clip}\left(\hat{r}_{\theta,\eta}(f(t_{k})|s(t_{k})),1-\epsilon,1+\epsilon\right)\hat{A}_{\eta}(s(t_{k}),f(t_{k}))\right\},$
where the dataset $\mathcal{D}_{\eta}^{T}$ contains $T$ days of simulation data generated under policy $\pi_{\eta}$ (including the atomic states, atomic actions and order of patients within the simulation of sequential making process for each day), and the vector $\hat{\textbf{A}}_{\eta}(\mathcal{D}_{\eta}^{T})=\left\{\left(\hat{A}_{\eta}(s(t_{k}),f(t_{k}))\right)_{k=0}^{m-1}\right\}_{t=1}^{T}$ consists of the estimated advantage functions for each decision epoch within the $T$ days using the collected data $\mathcal{D}_{\eta}^{T}$. The detailed steps of the PPO algorithm are given in Algorithm 1.
@@
1:for policy iteration r=0,1,2,,R1r=0,1,2,...,R-1 do
2:  Run policy $\pi_{\eta}$ for $T$ days and collect simulation dataset $\mathcal{D}_{\eta}^{T}$.
@@
Computing $\pi(f\mid s)$ in (5) is computational expensive due to the required summation. In implementation, we approximate it using the product $\prod_{n=1}^{q(t_{k})}\kappa\big(a^{n}(t_{k})\mid s^{n}(t_{k}),\sigma^{n}(t_{k})\big)$, evaluated along the realized processing order in simulation; correspondingly, we compute an approximate policy ratio $\hat{r}_{\theta,\eta}$. In Section 5.3, we further introduce a batching scheme that eliminates order dependence and yields a simpler, more direct mapping between the system-level policy $\pi$ and the atomic-level distributions $\kappa$.
We conclude this section with two important remarks. First, our atomic policy parameterization differs from the conventional ones used in PPO framework, which directly parameterize a randomized policy over the entire system-level action space and would otherwise require an intractably large policy network. Second, performance improvement in PPO relies on conservative policy updates (Proposition 1 in Appendix 10), which is infeasible under the deterministic policy setup. In particular, while atomic actions alone reduce the action space, combining them with traditional value-based ADP leads to severe policy “chattering.” This phenomenon arises because, when value functions are approximated rather than evaluated exactly, greedy policy updates may overreact to estimation errors, causing oscillations between improving and worsening policies (Bertsekas and others 2011). PPO’s proximal updates mitigate this issue. In other words, the combination of atomic actions and PPO is critical; see empirical results in Section 6.
5 Tailoring Design for Periodicity in MDP
@@ 5.1 Policy Network Design
Figure 2: Three policy network structures. The fully-connected network takes epoch index as part of the input and uses a single network to learn actions across all epochs. The fully-separate network trains $m$ separate networks for each of the $m$ epochs. The partially-shared network excludes the epoch index from input but uses different blocks of output neurons for actions of different epochs.
@@
    where $g_{i,j}$ is one of the $J\times J$ output neurons of the fully-connected network. This design uses the same set of parameters to generate outputs for different epochs, which captures temporal differences by incorporating time as part of the input.
@@
Combining advantages. The fully-connected and fully-separate designs each have their pros and cons, while the partially-shared policy network combines their strengths. The fully-connected network design uses the same set of parameters across all time epochs, naturally capturing similarities between epochs and enabling training with data from all time epochs. However, since the policy differences between epochs can be highly complex, a more sophisticated network (e.g., with additional hidden layers) may be needed to capture these nuances and hence more training data. The fully-separate network design accounts for differences between epochs with independent networks, but it would also require more time and data to train, as each network is only trained using data from its corresponding epoch. The partially-shared network design strikes a balance: shared parameters in the early layers capture commonalities between epochs for efficient training, while the separated last layer adapts to epoch-specific policy outputs. In the case study of a ten-pool system, the partially-shared design requires only a few thousand simulation data samples for training, yet outperforms the other two designs that need 100 times more data; see more in Section 6.4.
@@ 5.2 Policy Evaluation: Linear Value Function Approximation
In the standard PPO implementation, a separate neural network, known as the value network, is used alongside the policy network to approximate value functions. The value network is trained with targets estimated via Monte Carlo evaluation, which can have high variance in long-run average settings. In this paper, we adopt a linear value function approximation and leverage the queueing system structure to design proper basis functions. To train the value approximation, we use the Least-Squares Temporal Difference (LSTD) method proposed by Bradtke and Barto (1996), which provides more stable estimates in tabular and linear function approximation settings, as shown in Boyan (1999) and Lagoudakis and Parr (2003).
@@
Specifically, under the randomized atomic policy $\pi$, we decompose the original $J$-pool system into $J$ single-pool systems, and the relative value function $\hat{v}_{\pi,j}$ for the decomposed pool-$j$ system is solved via the post-action Poisson equation
@@
for a given post-action state $s_{j}$ (the superscript ++ dropped for simplicity). Here, $C_{j}q_{j}$ is the post-action holding cost with $q_{j}=\max(s_{j}-N_{j},0)$, $\hat{\gamma}_{\pi,j}$ is the average cost of the decomposed pool-$j$ system, and the cost-to-go term consists of: (a) the expected overflow cost $\mathbb{E}[\sum_{k=1}^{J}B_{j,k}\bar{f}_{j,k}]$, where $\bar{f}_{j,k}$ approximates the overflow assignments from class $j$ to other pools, and (b) the expected value function at the next post-action state $\mathbb{E}\left[\hat{v}_{\pi,j}(s’_{j})\right]$. To evaluate (a) and (b), we need to properly adjust the queueing dynamics for each decomposed pool-$j$ system by incorporating overflow assignments from other classes while excluding overflow assignments from class $j$ to other pools (treated as “diversions”). To achieve this, we approximate the probability of assigning a class-$j$ customer to pool $k$ under policy $\pi$ using $\bar{\kappa}_{j}(k|s_{j},j)$, which depends only on pool-$j$’s state information. Given the arrivals are Poisson and leveraging the Poisson thinning property, the outflows $\bar{f}_{j,k}$ and inflows $\bar{f}_{i,j}$ can be approximated as Poisson distributions with rates $\lambda_{j}\bar{\kappa}_{j}(k|s_{j},j)$ and $\lambda_{i}\bar{\kappa}_{i}(j|s_{j},i)$, respectively. Details on computing $\bar{\kappa}_{j}$ and deriving $\hat{v}_{\pi,j}$ are provided in Appendix 11.